Use a system of two equations in two variables to solve the problem.
At a store, ice cream cones cost $1.80 and sundaes cost $3.30. One day a total of 158 cones and sundaes were sold, and the receipts totaled $389.40. How many cones were sold?



Answer by ankor@dixie-net.com(22740)

Use a system of two equations in two variables to solve the problem.
At a store, ice cream cones cost $1.80 and sundaes cost $3.30.
One day a total of 158 cones and sundaes were sold, and the receipts totaled $389.40.
How many cones were sold?
:
let c = no. of cones
let s = no. of sundaes
:
c + s = 158
1.8c + 3.3s = 389.40
multiply the 1st equation by 3.3, subtract the 2nd equation
3.3c + 3.3s = 512.40
1.8c + 3.3s = 389.40
----------------------Subtraction eliminates s, find c
1.5c = 123.00
c = 123/1.5
c = 82 cones sold
:
You can confirm this yourself, there are 158 - 82 = 76 sundaes
see if the receipt total is correct with these values